# Modified Schwarzschild spacetime

### v1

@@ 3 Physical properties
and this integral diverges. This indicates only a coordinate type of singularity because the derivative of the radial coordinate with respect to the parameter of the null geodesic is regular on the horizon, and the same is true also for mass particles radially approaching the horizon, where the derivative is with respect to their proper time. This is the same behavior as with the Schwarzschild metric without extra dimensions. In the nontrivial case with the spacetime metric (2.11) for the considered ingoing signal, the amount of coordinate time needed to reach the horizon is given by
[…]
where the last expression is the result in the leading order of the approximation of small $\Delta/|a|$ for $n>0$. We can see that this result is finite which indicates that singularity of the $g_{00}$ component of the metric is not caused by an inappropriate choice of the time coordinate, but rather it suggests a physical singularity, already revealed by the Kretschmann scalar.
Another approach to characterization of properties of spacetimes is based on Landau–Lifshitz stress-energy pseudotensor [51] providing definition of conserved quantities. In a $D$-dimensional spacetime a conserved $D$-momentum $P^{\mu}$ associated with a $D-1$ dimensional space-volume region $\Omega$ can be calculated as
[…]
where $d\Sigma_{\nu}$ denotes $D-2$ dimensional oriented area element of the boundary of $\Omega$, and through object $h^{\mu\nu\lambda}$ we can obtain the total stress-energy pseudotensor as $t_{\textrm{tot}}^{\mu\nu}=h^{\mu\nu\lambda}_{\phantom{\mu\nu\lambda},\lambda}$. It is defined as
@@
$h^{\mu\nu\lambda}=\dfrac{1}{16\pi\kappa}\left[(-g)\left(g^{\mu\nu}g^{\lambda\sigma}-g^{\mu\lambda}g^{\nu\sigma}\right)\right]_{,\sigma}.$ (3.8)
For evaluation of the conserved quantities defined above for spacetimes with metrics (2.10) and (2.11), we chose the $3+n$ dimensional volume $\Omega$ as a sphere in the three-dimensional space part, so that the topology of its boundary $\partial\Omega$ is $S^{2}\times S^{1}\times...\times S^{1}$. The area element is given by $d\Sigma_{i}=dS_{i}$ and $d\Sigma_{A}=0$, where $dS_{i}$ corresponds to the area element of the sphere boundary. The only nonzero component of the $(4+n)$-momentum is energy $\mathcal{E}=P^{0}$, and it can be calculated as
[…]
To evaluate this integral, it is useful to replace the radial coordinate $r$ with coordinate $\rho$ defined by the transformation $r=\rho\left(1-a/(4\rho)\right)^{2}$. This new coordinate is usually called isotropic radial coordinate because in the case with $1+3$ dimensional Schwarzschild metric, the space part of the metric becomes conformally Euclidean. In our case, this is true only for the three-dimensional and $n$-dimensional sections separately. For positive values of $a$, the isotropic radial coordinate $\rho$ cowers the entire spacetime except for the central singularity twice, while for negative $a$, it cowers twice only the region above the horizon. To avoid the double coverage, we may consider $\rho$ only in the interval $[1/4,\infty)$.
In the trivial case (2.10) we have
where functions $\chi_{+}$ and $\chi_{-}$ are defined as $\chi_{\pm}=1\pm a/(4\rho)$, and space coordinates $x^{i}$ are Cartesian such that $(x^{1})^{2}+(x^{2})^{2}+(x^{3})^{2}=\rho^{2}$. Again, this is only a trivial extension of the Schwarzschild solution, now written in isotropic coordinates. The conserved energy calculated through formula (3.9) for boundary $\partial\Omega$ chosen in the aforementioned way, at some fixed value of the radial coordinate $\rho$, is
@@
$\mathcal{E}(\rho)=-\dfrac{a}{2\kappa}\left(\chi_{-}\right)^{7},$ (3.11)
and in the limit of $\Omega$ covering entire spacetime, $\rho\to\infty$, we have $\mathcal{E}=-a/(2\kappa)=M$. This is the same result as in the case with the Schwarzschild spacetime without additional space dimensions.
In the nontrivial case (2.11), the spacetime metric can be rewritten as
$ds^{2}=-\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{-2\frac{n-1}{n+1}}dt^{2}+\left(\chi_{-}\right)^{4}\delta_{ij}dx^{i}dx^{j}+\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{\frac{2}{n+1}}\delta_{AB}d\zeta^{A}d\zeta^{B}.$ (3.12)
The formula (3.9) then implies
$\mathcal{E}(\rho)=\dfrac{a\left(\chi_{+}\right)^{\frac{4n}{n+1}}\left(\chi_{-}\right)^{4\frac{n+2}{n+1}}}{4(n+1)\kappa}\left(\dfrac{n}{\chi_{+}}-\dfrac{n+2}{\chi_{-}}\right),$ (3.13)
and for the entire spacetime we have $\mathcal{E}=-a/(2(n+1)\kappa)$. We may also notice that for $n=0$, we retrieve the result of the trivial case (3.11).
In summary, we have $a=-2\kappa M$, and $\mathcal{E}=M$ for the Schwarzschild spacetime as well as for its trivial extensions (2.10). On the other hand, in the nontrivial case
$a=2\dfrac{n+1}{n-1}\kappa M,\quad\mathcal{E}=-\dfrac{1}{n-1}M,$ (3.14)
and there is also horizon singularity at $r=-a$ if $a$ is negative and $n>1$. For $n=0$, both cases coincide with the Schwarzschild spacetime, and in the nontrivial case with $n=1$, the mass given by the Newtonian limit is zero, $M=0$, and constant $a$ cannot be fixed in the usual way. These unusual properties set the nontrivial case (2.11) apart from the trivial one (2.10).
@@ 4 Discussion
There are three distinctive classes of Ricci-flat spacetimes with metric given by the ansatz (2.3). For $f=1$, we have $1+3+n$ dimensional Minkowski spacetime. We focused on more intriguing cases given by metrics (2.10) and (2.11). The first is a direct product of the $1+3$ dimensional Schwarzschild spacetime and $n$-dimensional Euclidean space. If we assume microscopic extra dimensions with $n$-torus topology, such spacetimes do not differ from the ordinary Schwarzschild spacetime from the point of view of physical objects larger than the size of extra dimensions. Most interesting Ricci-flat solutions are given by the metric (2.11), and in the previous section, we found their nonstandard properties.
Comparing them with other cases mentioned above is helpful in appreciating their unusual features. This can be summarized in table 1.
horizons singularities Newt. lim. mass cons. energy 1. Minkowski ($a=0$) none none $M=0$ $\mathcal{E}=0$ 2. trivial ($a>0$) none $r=0$ $M=\frac{-a}{2\kappa}<0$ $\mathcal{E}=M<0$ 3. trivial ($a<0$) $r=-a$ $r=0$ $M=\frac{-a}{2\kappa}>0$ $\mathcal{E}=M>0$ 4. nontrivial ($a>0$) none $r=0$ $M=\frac{n-1}{n+1}\frac{a}{2\kappa}>0$ $\mathcal{E}=\frac{-a}{2(n+1)\kappa}<0$ 5. nontrivial ($a<0$) $r=-a$ $r=\{0,-a\}$ $M=\frac{n-1}{n+1}\frac{a}{2\kappa}<0$ $\mathcal{E}=\frac{-a}{2(n+1)\kappa}>0$
Table 1: Summary of properties of all possible Ricci-flat spacetimes with the metric of the form (2.3), including Minkowski spacetime and solutions (2.10) and (2.11). Note that in case 4., the existence of the horizon singularity at $r=-a$ occurs only for the number of extra dimensions $n>1$.
As we can see, only cases 1. and 3 are well-behaved and have the same properties as Minkowski and Schwarzschild spacetimes. Case 2. is nonphysical because it requires negative mass defined through the Newtonian limit and conserved energy from the Landau–Lifshitz pseudotensor, which also coincide with each other, $\mathcal{E}=M$. It also lacks a horizon, and therefore, it has a naked singularity. These reasons also exclude the possibility of negative values of constant $a$ for the Schwarzschild solution with no extra dimensions.
Cases 4. and 5. with spacetime metric of the form (2.11) show some peculiar properties. The sign of constant $a$ determines the existence of a horizon. It is located at $r=-a$, if $a$ is negative. There is a central singularity at $r=0$ in both cases, as indicated by the Kretschmann scalar (3.3), but there is also a horizon singularity in the case with negative $a$ and $n>1$. In case 4., there is a naked singularity, attractive gravitational effect in the Newtonian limit, $M>0$; However, the conserved energy $\mathcal{E}$ is negative. In case 5., there is a horizon singularity, the conserved energy is positive, but gravity acts repulsively in the Newtonian limit.
Another important difference between cases 2. and 3. with the spacetime metric (2.10) and cases 4. and 5. with (2.11) is the size of extra dimensions, if they are compactified with the $n$-torus topology. With the trivial extension of the Schwarzschild spacetime, we have $g_{AB}=\delta_{AB}$. Hence, the size of extra dimensions remains the same throughout the entire spacetime. However, we have $g_{AB}=\left(1+a/r\right)^{2/(n+1)}\delta_{AB}$ in the nontrivial case. In the case with positive $a$, there is no horizon, and the size of extra dimensions smoothly transits from a constant in large distances from the center to infinite divergence at the central singularity. This behavior at the center and far away is the same for the case with negative $a$, but in this case, the size of extra dimensions also drops to zero at the horizon. As we have shown, this horizon corresponds to a physical singularity.
Solutions with the nontrivial extension of the Schwarzschild spacetime (2.11) may, in principle, correspond to exterior solutions around some spherically symmetric objects [52]. However, such objects would have nonphysical properties. If we demand the conserved energy $\mathcal{E}$ to be positive, such an object would manifest repulsive gravity. A system containing such an exotic object and another massive astrophysical object could then accelerate itself, which is physically nonsensical. To avoid this problem, the conserved energy $\mathcal{E}$ must be negative, which is also problematic. This suggests that the solution (2.11) is nonphysical.
@@
In addition to problems with nonphysical properties, there are potential problems with instability. Some higher-dimensional extensions of the Schwarzschild spacetime tend to be unstable [53], which also includes brane solutions [54, 55]. Case 4. in the table contains naked singularity, which also hints at potential problems with stability [56, 57].
Finally, let us point out that only case 3. in the table can be maximally extended in the same way as the Schwarzschild spacetime [58]. In cases 2., 4. and 5., either the horizon singularity or lack of the horizon prevents the maximal extension by passing to Kruskal–Szekeres coordinates.
This paper has provided an analysis of possible extensions of the Schwarzschild spacetime with $n$ extra dimension and $SO(3)\times E(n)$ symmetry. We may conclude that only solutions with the spacetime metric (2.10) and choice of negative values for constant $a$ are physically well-behaved. They represent a trivial extension of the $1+3$ dimensional Schwarzschild spacetime. We have also found more intriguing solutions (2.11). Despite their problematic features, they provide an interesting testing ground for general relativity since they represent one of not many examples of vacuum solutions of Einstein field equations. Problems with some of their nonphysical properties may perhaps be avoided by generalizing and modifying them further, but we leave this for further work.

### v2

@@ 3 Physical properties
and the Newtonian mass is related to the constant $a$ through a different relation for the Newtonian potential
$\phi=-\dfrac{n-1}{n+1}\dfrac{a}{2}\dfrac{1}{\rho}=-\kappa M_{\textrm{New.}}\dfrac{1}{\rho}.$ (3.6)
A consequence of a negative value of $a$ is the existence of a horizon. However, in the modified case with the spacetime metric (2.11), the sign in the relation above is opposite, which implies naked singularities for positive masses $M_{\textrm{New.}}$. Negative $a$ and the existence of a horizon is then associated with negative mass $M_{\textrm{New.}}$, which means repulsive gravity in the Newtonian limit. Other examples of different types of black holes with such a property are studied, for example, in [52]. We may also notice that the presence of extra dimensions does not change the position of the horizon or its properties considerably. For example, a contrary case can be found in [53]. Note also that for $n=1$, the spacetime metric in the non-trivial case is (2.2), and the Newtonian potential vanishes, which means zero Newtonian mass $M_{\textrm{New.}}$.
The physical length of periodicity in extra dimensions is proportional to $p_{\textrm{phys.}}\propto\sqrt{g_{AB}}$. We can see that the size of extra dimensions is constant in entire spacetime in the trivial case (2.10), while in the non-trivial case (2.11), it depends on the radial coordinate, collapsing to zero size on the horizon if $a$ is negative and also diverging to infinity at the central singularity regardless of the sign of $a$. Let us assume that the maximal size of extra dimensions is at the Planck distance $l_{\textrm{Pl}}$ from the center, since one cannot approach any point at a closer distance. By calculating the physical radial distance as $r_{\textrm{phys.}}=\int_{0}^{r}\sqrt{g_{rr}}dr$ and taking the limit $r\ll a$, we find the ratio of physical periodicity at the Planck distance from the center to its value very far from the origin,
$\lambda=\dfrac{p_{\textrm{phys.}}(r_{\textrm{phys.}}=l_{\textrm{Pl}})}{p_{\textrm{phys.}}(r_{\textrm{phys.}}\to\infty)}\approx\left[1.22\cdot 10^{38}\dfrac{n+1}{n-1}\dfrac{M_{\textrm{New.}}}{M_{\textrm{sol.}}}\right]^{\frac{2}{3}\frac{1}{n+1}},$ (3.7)
where we have used also (3.6), and $M_{\textrm{sol.}}$ denotes the solar mass. For supermassive black holes with a mass of the order $10^{10}M_{\textrm{sol.}}$ and the minimal number of extra dimensions allowing horizon singularity $n=2$, we have $\lambda\approx 6\cdot 10^{10}$, and this factor becomes smaller for smaller masses as well as for higher numbers of extra dimensions. This means that if the size of extra dimensions is microscopic far from the origin, it remains very small even near the central singularity.
In addition to the Newtonian mass, there are also other ways to assign a mass-like quantity to spacetimes or spacetime regions. Komar mass [54] associated with space region $\Omega$ can be calculated as [55, 56]
$M_{\textrm{Kom.}}=-\dfrac{1}{8\pi\kappa}\int\limits_{\partial\Omega}\star d\sigma,$ (3.8)
where $\star d\sigma$ is dual of two form $d\sigma$ defined through timelike Killing one form $\sigma$. In the case with spacetime metric of the form (2.10) as well as (2.11), we have $\sigma=g_{00}dt$, and $\star d\sigma=g_{00,r}d\vartheta\wedge d\varphi\wedge d\zeta^{1}\wedge...\wedge d\zeta^{n}$. By choosing the volume region $\Omega$ as a sphere in the three-dimensional space part with radius $r=R$, so that the topology of its boundary $\partial\Omega$ is $S^{2}\times S^{1}\times...\times S^{1}$, we find
$M_{\textrm{Kom.}}(R)=-\dfrac{1}{2\kappa}R^{2}g_{00,r}(R)L^{n},$ (3.9)
where $L^{n}=\int d\zeta^{1}...d\zeta^{n}$ is $n$-dimensional volume of the extra dimensions far from the origin, $r\to\infty$, where $g_{AB}\to\delta_{AB}$. By taking the limit $R\to\infty$, we find that the Komar mass coincides with the Newtonian mass discussed above up to the $n$-dimensional volume factor, $M_{\textrm{Kom.}}=M_{\textrm{New.}}L^{n}$.
Another approach to defining conserved quantities characterizing properties of spacetimes is based on pseudotensors. Here, we consider Einstein [57] and Landau–Lifshitz [58] stress-energy pseudotensors. Mass defined through the Einstein stress-energy pseudotensor can be calculated as [59]
$M_{\textrm{Ein.}}=\int\limits_{\partial\Omega}E_{0}^{\phantom{0}0i}dS_{i},$ (3.10)
@@
$E_{\mu}^{\phantom{\mu}\nu\rho}=\dfrac{1}{16\pi\kappa}\dfrac{g_{\mu\sigma}}{\sqrt{-g}}\left[(-g)\left(g^{\nu\sigma}g^{\rho\lambda}-g^{\rho\sigma}g^{\nu\lambda}\right)\right]_{,\lambda}.$
Similarly, the mass defined through Landau–Lifshitz stress-energy pseudotensor is given by
$M_{\textrm{L.,L.}}=\int\limits_{\partial\Omega}L^{00i}dS_{i}$ (3.11)
$L^{\mu\nu\rho}=\dfrac{1}{16\pi\kappa}\left[(-g)\left(g^{\mu\nu}g^{\rho\sigma}-g^{\mu\rho}g^{\nu\sigma}\right)\right]_{,\sigma}.$
By considering the same region $\partial\Omega$ as previously and by taking the limit $R\to\infty$, we find that $M_{\textrm{Ein.}}=M_{\textrm{L.,L.}}=-aL^{n}/(2\kappa)$ for the trivial case (2.10) with the use of its form in Newtonian limit (3.4). These two masses also equal the Komar mass as well as the Newtonian mass up to the factor $L^{n}$. On the other hand, in the non-trivial case with the metric of the form (3.5), we obtain
$M_{\textrm{Ein.}}=M_{\textrm{L.,L.}}=-\dfrac{1}{n+1}\dfrac{a}{2\kappa}L^{n}.$ (3.12)
Einstein and Landau–Lifshitz masses coincide, but they differ from Komar and Newtonian masses in this case.
Finally, the ADM mass [60, 61, 62] can be calculated as [63]
$E=\dfrac{1}{16\pi\kappa}\oint\limits_{\partial\Omega}\sqrt{\gamma}\gamma^{\mu\nu}\gamma^{\alpha i}\left(\gamma_{\mu\alpha,\nu}-\gamma_{\mu\nu,\alpha}\right)dS_{i},$ (3.13)
where $\gamma_{\mu\nu}$ is metric of chosen spacelike hypersurface $\partial\Omega$. We choose it as in the previous cases. For the Schwarzschild spacetime with no extra dimensions, we have $E(R)=-(R/\kappa)\sqrt{1+a/R}$, and to obtain a meaningful result, we have to subtract the same quantity evaluated in the Minkowski spacetime with $a$ set to zero, $E_{0}(R)=-R/\kappa$. The total ADM mass then can be found as $M_{\textrm{ADM}}=\lim\limits_{R\to\infty}\left(E(R)-E_{0}(R)\right)=-a/(2\kappa)$. In the trivial case with extra compact dimensions (2.10), we have an expected result,
@@
$M_{\textrm{ADM}}=\lim\limits_{R\to\infty}\left(E(R)-E_{0}(R)\right)=\lim\limits_{R\to\infty}\dfrac{R}{\kappa}\left(1-\sqrt{1+\dfrac{a}{R}}\right)L^{n}=-\dfrac{a}{2\kappa}L^{n},$ (3.14)
while in the non-trivial case with (2.11), we find
$M_{\textrm{ADM}}$ $=$ $\lim\limits_{R\to\infty}\left(E(R)-E_{0}(R)\right)=$
$=$ $\lim\limits_{R\to\infty}\dfrac{R}{\kappa}\left[1-\left(1+\dfrac{1}{2}\dfrac{n+2}{n+1}\dfrac{a}{R}\right)\left(1+\dfrac{a}{R}\right)^{\dfrac{1}{2}\dfrac{n-1}{n+1}}\right]L^{n}=-\dfrac{2n+1}{n+1}\dfrac{a}{2\kappa}L^{n}.$
The ADM mass is the same as other masses in the trivial case, while it differs from all of them in the non-trivial case.
All masses calculated for the non-trivial case reduce to results for the trivial case by setting $n=0$. Therefore, all results can be summarized as
$M_{\textrm{New.}}L^{n}=M_{\textrm{Kom.}}=\dfrac{n-1}{n+1}\dfrac{a}{2\kappa}L^{n},$ (3.16)
$M_{\textrm{Ein.}}=M_{\textrm{L.,L.}}=-\dfrac{1}{n+1}\dfrac{a}{2\kappa}L^{n},\quad M_{\textrm{ADM}}=-\dfrac{2n+1}{n+1}\dfrac{a}{2\kappa}L^{n},$
with $n=0$ corresponding both to the original Schwarzschild spacetime as well as to its trivial extension with an arbitrary number of extra dimensions (2.10). We can see that signs of $M_{\textrm{Ein.}}$, $M_{\textrm{L.,L.}}$ and $M_{\textrm{ADM}}$ are opposite to signs of $M_{\textrm{New.}}$ and $M_{\textrm{Kom.}}$ for $n\geq 2$. Negative ADM mass of Kaluza–Klein bubbles has been found, for example, in [64].
@@ 4 Discussion
This paper has provided an analysis of possible extensions of the Schwarzschild spacetime with $n$ extra dimensions and $SO(3)\times E(n)$ symmetry. There are three distinctive classes of static vacuum solutions of the Einstein field equations with metric given by the ansatz (2.3). For $f=1$, we have $1+3+n$ dimensional Minkowski spacetime, and more interesting cases are given by metrics (2.10) and (2.11). The first is a direct product of the $1+3$ dimensional Schwarzschild spacetime and $n$-dimensional Euclidean space. If we assume microscopic extra dimensions with $n$-torus topology, such spacetimes do not differ from the ordinary Schwarzschild spacetime from the point of view of physical objects larger than the size of extra dimensions. Most intriguing solutions are given by the metric (2.11), and in the previous section, we investigated their nonstandard properties.
For the number of extra dimensions $n\geq 2$, the Kretschmann scalar (3.1) of the spacetime metric (2.11) diverges on the horizon, r=ar=-a, which indicates a physical singularity on the horizon. Extra dimensions collapse on this horizon, which is a feature of Kaluza–Klein bubbles. There are three different masses in this case (3.16). Newtonian mass and Komar mass coincide, $M_{\textrm{Kom.}}=M_{\textrm{New.}}L^{n}$, and Einstein, Landau–Lifshitz and ADM masses are given by
$M_{\textrm{Ein.}}=M_{\textrm{L.,L.}}=-\dfrac{1}{n-1}M_{\textrm{Kom.}},\quad M_{\textrm{ADM}}=-\dfrac{2n+1}{n-1}M_{\textrm{Kom.}}.$ (4.1)
Existence of a physical singularity on the horizon is then associated with positive Einstein, Landau–Lifshitz, and ADM masses, and negative Newtonian and Komar masses. Three different masses originate from three independent components of the metric (2.11). The Newtonian and Komar masses are given by only $g_{00}$ component, Einstein and Landau–Lifshitz masses have been calculated through the determinant (g)(-g), $g_{00}$ and $g_{ij}$ components, and to calculate the ADM mass we have used $g_{rr}$ and $g_{AB}$ components. A similar example is a class of spherically symmetric spacetimes with three space dimensions and with only two independent components that yield only two independent masses [65]. In the trivial case (2.10) with a positive constant $a$, there is no horizon, and all masses discussed above are equal, and they are negative.
@@
Since different masses (3.16) have different signs indicating nonphysical properties, we should also expect problems with instability. Some higher-dimensional extensions of the Schwarzschild spacetime tend to be unstable [66], which also includes Kaluza–Klein bubbles [36] and brane solutions [67, 68]. Cases with a positive constant $a$ feature a naked singularity, which also hints at potential problems with stability [69, 70]. The spacetime metric (2.11) then represents a case in which different masses are associated with problematic properties.